A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. She knows that Solution A is 65% salt and Solution B is 90% salt. She wants to obtain 70 ounces of a mixture that is 80% salt. How many ounces of each solution should she use?
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. She knows that Solution A is 65% salt and Solution B is 90% salt. She wants to obtain 70 ounces of a mixture that is 80% salt. How many ounces of each solution should she use?
------------
Assuming there is some compound that can be 90%, or even 65%, in some solvent, which I doubt:
"Salt" is not necessarily NaCl. There are a lot of salts.
---------------
A + B = 70 --- total solution
65A + 90B = 70*80 ---- total salt
---------
etc

Answer by ikleyn(52787)

.

The salt solution CAN NOT have concentration higher than 28%.

After that the solution becomes SATURATED.


To know this fact from  SCIENCE  is  MUCH  MORE  IMPORTANT  thing than to "solve" mechanically problems,
posed incorrectly,  without understanding the subject.


See this article from Wikipedia,  the free encyclopedia

https://en.wikipedia.org/wiki/Saline_watet.


Happy learning (!)


Penalty to the author of this  FAKE  problem.